You have at least $30 in change in your drawer, consisting of
dimes and quarters. Write an inequality that shows the different number of coins in your drawer.

Answer by venugopalramana(3286)

You have at least $30 in change in your drawer, consisting of
dimes and quarters. Write an inequality that shows the different number of coins in your drawer.
LET NUMBER OF DIMES =D...THEIR VALUE =10*D=10D CENTS
LET NUMBER OF QUARTERS = Q...THEIR VALUE =25*Q=25Q CENTS.
TOTAL VALUE =10D+25Q
HENCE
10D+25Q>=30*100=3000 CENTS
2D+5Q>=600